OpenStudy (anonymous):

Bob has $2.90 in pennies, nickels and dimes. He has the same number of pennies and nickels. If the number of dimes is 5 more than the number of nickels, how many of each type of oin does he have?

OpenStudy (anonymous):

x+x+x+5=2.90 3x=2.1 x=7 7 pennies 7 nickel and 10 dimes

OpenStudy (anonymous):

oh cool, thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

wait i think its wrong

OpenStudy (dumbcow):

10D +5N +P = 290 P=N D = 5+N use substitution 10(5+N) +5N +N = 290 50 +10N +5N +N = 290 16N+50 = 290 16N = 240 N = 15 then P = 15 D = 5+15 = 20